OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 72°?

OpenStudy (cwrw238):

interior angle = 180 - 72 = 108 exterior angle was found by x = 360 / n and x is 72 so you can calculate no. of sides n and sum of angles will be 108*n

OpenStudy (anonymous):

that's what i thought! thank you

OpenStudy (cwrw238):

yw